OpenStudy (anonymous):

If you were to use the substitution method to solve the following system, choose the new system of equations that would result if x was isolated in the second equation.

hero (hero):

Where's the first and second equations?

OpenStudy (anonymous):

3x + y + 2z = 1 x + 2y – z = –8 3x – y – 3z = 18

hero (hero):

Suggestion: Do not use substitution method with system of three equations. Use matrices.

OpenStudy (anonymous):

i dont even know what that means math is sooo confusing

hero (hero):

Yeah, and for you, I can see why. They want to torture you by making you use substitution method rather than matrices.